OpenStudy (anonymous):

Find the equation of the tangent line to the hyperbola at the given point: (x^2/a^2) - (y^2/b^2) = 1 (m, n) and write your answer in the form y=f(x) for some function f(x).

OpenStudy (anonymous):

the point \((m,n)\)?

OpenStudy (anonymous):

I used implicit diff. to get y' = (2a^2y)/(2b^2x) but then what's next?

OpenStudy (anonymous):

Yes the point (m, n)

OpenStudy (anonymous):

cancel the twos

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

replace \(x\) by \(m\) and \(y\) my \(n\) for your slope

OpenStudy (anonymous):

so a^2y / b^2x becomes a^2n / b^2m ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then use the point - slope formula

OpenStudy (anonymous):

So, y = (a^2n / b^2m)(x-m) + n ?

OpenStudy (anonymous):

yeah, nice and ugly

OpenStudy (anonymous):

i actually didn't check your derivative but i assume it is right

OpenStudy (anonymous):

hmm maybe it is not

OpenStudy (anonymous):

Yeah its not... >.<

OpenStudy (anonymous):

shouldn't it be \[\frac{b^2x}{a^2y}\]?

OpenStudy (anonymous):

I still get the wrong answer with that

OpenStudy (anonymous):

you still need to replace m and n for x and y

OpenStudy (anonymous):

\[\frac{b^2m}{b^2n}(x-m)+n\] try that

OpenStudy (freckles):

put an a^2 on bottom there instead of that b^2 :p

OpenStudy (anonymous):

it is late

OpenStudy (freckles):

yes I know it is time for the old folks to retire

OpenStudy (freckles):

that includes me of course

OpenStudy (anonymous):

((b^2m) / (a^2n))(x-m) + n thats right

OpenStudy (anonymous):

:D Thanks so much!! @satellite73 @freckles :D

OpenStudy (anonymous):

yw

OpenStudy (freckles):

I'm just satellite's cheerleader on this one! Go satellite!!!

OpenStudy (freckles):

And Tracy! :)

OpenStudy (anonymous):

Ahaha XD thats an important role @freckles !!